Three genes are located on the same chromosome and are known to be 8 mu apart between A and B, and 12 mu apart between B and C. The heterozygous AaBbCc genotype individuals were crossed to homozygous recessive aabbcc genotype individuals (test cross) and observed 3 double crossover offsprings. If this cross produces a total of 500 offsprings, what are the expected numbers of double crossover offsprings and calculate the Interference (1) value. (3 points) 4·
Answer #1

Answer:

Expected double crossover offsprings = 5

Interference = 0.4

Explanation:

A----8mu----B----12mu----C

Distance between genes (mu) = Recombination frequency (%)

Expected double crossover frequency = (RF between A&B) (RF between B&C)

= 8% x 12%

= 0.08 x 0.12

= 0096

Number of double crossover offsprings = 0.0096*500 = 4.8 or 5

Coefficient of Coincidence (COC) = Observed double crossovers / Expected double crossovers

= 3/5 = 0.6

Interference (I)= 1 – COC

I = 1- 0.6 = 0.4